OpenStudy (anonymous):

Help Solve? Please give steps. 3d + d - 7 = 25/4

OpenStudy (abdullah_me):

3d + d - 7 = 25/4 4d-7=25/4 4d=25/4 + 7 d=25+ 4*7 d=25+28=53